### v1

@@ 7 Conclusion and future outlook
Figure SM4: Numerical test problems for MGRIT applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM5). Left column: The wave-speed function $\alpha(x,t)$. Right column: Space-time contour of the corresponding solution $e(x,t)$ of the PDE (cons-lin). From top to bottom, the wave-speed functions are: $\alpha(x)=1$, $\alpha(x)=\frac{1}{2}\big{(}1+\cos^{2}(\pi x)\big{)}$, $\alpha(x,t)=-\sin^{2}\big{(}\pi(x-t)\big{)}$, and $\alpha(x,t)=\cos(2\pi x)\cos(2\pi t)$, respectively.
In this section, we show results for MGRIT used to solve discretizations of the PDE (cons-lin) for four different wave-speed functions $\alpha$. The different wave-speed functions are shown in Fig. SM4, along with contours of the corresponding PDE solution in space-time. In all cases, the PDE initial condition is $e(x,0)=\sin^{4}(\pi x)$, the time domain is $t\in[0,4]$, and the (fine-grid) time-step is $\delta t=0.85h$.
@@
Figure SM5: Residual histories for MGRIT V-cycles applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM5). Left column: 1st-order accurate discretizations. Right column: 3rd-order accurate discretizations. Each of the four rows corresponds with the test problem shown in the corresponding row of Fig. SM4. Solid lines correspond to GLF numerical fluxes, and dashed lines correspond to LLF numerical fluxes. Note that for the constant wave-speed problem (top row), the GLF and LLF numerical fluxes are identical (see discussion in Section SM5.1).
@@
Figure SM6: Residual histories for MGRIT V-cycles applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM5). The test problems here are the same as the GLF problems considered in the first two rows of Fig. SM5, except that the coarse-grid operator does not include a truncation error correction. Note: All vertical axes are on the same scale ranging from $10^{-10}\to 10^{5}$, while those in Fig. SM5 range from $10^{-10}\to 1$.
MGRIT residual history plots are then shown in Fig. SM5. These results are for multilevel V-cycles, using a coarsening factor of $m=8$ on all levels, and coarsening is continued until fewer than two points in time would result. FCF-relaxation is used on all levels. Coarse-grid operators on deeper levels in the hierarchy have the same structure as the operator (SM64) on level two, with the truncation errors being computed recursively (see discussion in [14, Section 3.4]). All coarse-grid linear systems are approximately solved with GMRES, which is iterated until residual is reduced in norm by by 0.01 or until 10 iterations are reached. The parameter $\vartheta$ that specifies exactly where inside each time interval the wave-speed and dissipation coefficients is frozen is set to zero. The initial MGRIT iterate is a vector with uniformly random entries, and MGRIT is iterated until the 2-norm of the space-time residual is reduced by at least 10 orders of magnitude.
@@
For comparative purposes, we include in Fig. SM6 MGRIT residual histories for a subset of the same problems from Fig. SM5, except that no approximate truncation error correction has been used on coarse levels. That is, the coarse-grid operator used in the tests in Fig. SM6 is (SM64) with ${\cal T}_{\rm ideal}^{n}={\cal T}_{\rm direct}^{n}=0$, or simply the rediscretized semi-Lagrangian method, $\Psi^{n}={\cal S}^{n,m\delta t}_{p,1_{*}}$. Evidently, the approximate truncation error correction substantially improves the convergence rate of MGRIT, and appears necessary for the solver to converge in a number of iterations $k\ll n_{t}/(2m)$—the number of iterations at which FCF-relaxation sequentially propagates the initial condition across the entire time domain [21].
SM6 Approximating truncation error for linearized method-of-lines discretization
@@
In this section, we describe our approach for approximating the truncation error operator ${\cal T}_{\rm ideal}^{n}$ in the coarse-grid operator (58), in the context where the fine-grid operator is a linearized method-of-lines discretization. Recall that the particular details of the linearized problem are given in Section 4. This situation is distinct from that in Section SM5, which developed an approximation of ${\cal T}_{\rm ideal}^{n}$ in the context of the fine-grid operator being a standard, linear method-of-lines discretization. That being said, the heuristics we use in this section here are heavily inspired by the work in Section SM5 given the excellent results in Fig. SM5; we strongly recommend reading Section SM5 before the current section.
@@
As in Section SM5, we suppose that the approximate truncation error operator takes the form
${\cal T}_{\rm ideal}^{n}=\textrm{approximation}\bigg{(}\sum_{j=0}^{m-1}\Big{(}\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big{(}\delta t\widehat{L}^{n+j}\big{)}^{p+1}\Big{)}\bigg{)}.$ (SM76)
@@
Note that the truncation error operator in Section SM5 was written generally for a $q$th-order ERK method, but here we have fixed $q=p=2k-1$ since we only consider discretizations of the nonlinear PDE (cons) with equal spatial and temporal orders of accuracy.
@@
The rest of this section is organized as follows. Section SM6.1 presents an approximation for spatial discretization errors ^$\widehat{{\cal E}}^{n+j}$ in (SM76). Section SM6.2 presents an approximation for the powers of the spatial discretization $\big{(}\widehat{L}^{n+j}\big{)}^{p+1}$ appearing in (SM76). The final approximation for (SM76) is presented in Section SM6.3.
SM6.1 Approximating the spatial discretization error
@@
Estimating rigorously the error term $\widehat{E}^{n+j}$ in the present setting is quite complicated. So, we instead pursue a heuristic approach based the result in Lemma SM5.3. Recall that Lemma SM5.3 estimates the error of the LF numerical flux (12) when: (i) reconstructions of the wave-speed in the flux are replaced with exact evaluations of the wave-speed, and (ii) the linear reconstructions $e_{i+1/2}^{\pm}$ use standard weights. This differs from the present setting, in which the LF numerical flux (12) uses: (i) reconstructions of the wave-speed that are based on the current linearization point (see Section 4.2), and (ii) the linear reconstructions $e_{i+1/2}^{\pm}$ use non-standard weights (see Section 4.3). Nonetheless, we suppose that the dominant error term in the numerical flux takes the same form as it does in the case of the standard linear discretization described above, and as is analyzed in Lemma SM5.3 of Section SM5. That is, we suppose (dropping temporal superscripts for readability): [^12] It is possible to analyze the error $\widehat{E}_{i+1/2}$ more rigorously than this; however, we do not present such details here since they do not appear necessary or helpful for our purposes.
$\widehat{E}_{i+1/2}\approx\frac{\nu_{i+1/2}}{2}\big{(}e_{i+1/2}^{-}-e_{i+1/2}^{+}\big{)}.$ (SM77)
@@
Thus, our first step for developing an expression for (SM76) is to estimate the difference $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ in (SM77). Recall from Section 4.3 that the exact form of the linearized reconstructions $e_{i+1/2}^{\pm}$ depends on the order of the reconstructions $u_{i+1/2}^{\pm}$ in the outer nonlinear iteration and on the linearization scheme being used. Specifically, if 1st-order reconstructions are used then $e_{i+1/2}^{\pm}$ are just standard linear reconstructions. For higher-order reconstructions $u_{i+1/2}^{\pm}$, the reconstructions are nonlinear since they depend on $\bar{\bm{u}}$ (supposing WENO weights are used, that is). The linearized reconstructions $e_{i+1/2}^{\pm}$ are then either: (i) approximate gradients of $u_{i+1/2}^{\pm}$ if the approximate Newton linearization (53) is used; or (ii) weighted reconstructions using the same weights as were used to compute $u_{i+1/2}^{\pm}$ if the Picard linearization (54) is used.
@@
$e_{i+1/2}^{-}=\sum\limits_{\ell=0}^{k-1}b_{i}^{\ell}(\bar{\bm{u}})\big{(}R^{\ell}\bar{\bm{e}}\big{)}_{i},\quad e_{i+1/2}^{+}=\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\big{(}\widetilde{R}^{\ell}\bar{\bm{e}}\big{)}_{i+1}.$ (SM78)
Our starting point for estimating $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ is the following corollary of Lemma SM4.5, which considered estimates for the standard polynomial reconstructions $\big{(}\widetilde{R}^{\ell}\bar{\bm{e}}\big{)}_{i},\big{(}{R}^{\ell}\bar{\bm{e}}\big{)}_{i}$.
Corollary SM6.1 (Error estimate for weighted reconstructions using $12k-$ 1 cells).
@@
$\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i}^{\ell}\big{(}\widetilde{R}^{\ell}\bar{\bm{e}}\big{)}_{i}=e(x_{i-1/2})-\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[\widetilde{b}_{i}^{\ell}\,\widetilde{\zeta}^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i,\ell}(x_{i-1/2})}\right],$ (SM79)
@@
$\sum\limits_{\ell=0}^{k-1}{b}_{i}^{\ell}\big{(}{R}^{\ell}\bar{\bm{e}}\big{)}_{i}=e(x_{i+1/2})-\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[b_{i}^{\ell}\,\zeta^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i,\ell}(x_{i+1/2})}\right].$ (SM80)
Proof SM6.2.
The result follows by substituting the error estimates of $\big{(}\widetilde{R}^{\ell}\bar{\bm{e}}\big{)}_{i}$ and $\big{(}{R}^{\ell}\bar{\bm{e}}\big{)}_{i}$ given in Lemma SM4.5 and then using the consistency of the weights.
@@
$e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $=\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i+1,\ell}(x_{i+1/2})}-b_{i}^{\ell}\,\zeta^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{i,\ell}(x_{i+1/2})}\right],$ (SM81)
$\approx\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\Big{[}\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}-b_{i}^{\ell}\,\zeta^{\ell}\Big{]}\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg{|}_{\xi_{\ell}},$ (SM82)
@@
where $\{\xi_{\ell}\}$ are unknown points. The approximation made here is that, for fixed $\ell$, we can replace the two unknown points $\xi_{i,\ell}(x_{i+1/2})$ and $\xi_{i+1,\ell}(x_{i+1/2})$ by single unknown point $\xi_{\ell}$. This is true to at least lowest order because $\xi_{i,\ell}(x_{i+1/2})$ and $\xi_{i+1,\ell}(x_{i+1/2})$ are at most a distance of ${\cal O}(h)$ apart (see Lemma SM4.5 for why this is true).
@@
Recalling that ultimately we want to discretize $\widehat{E}_{i+1/2}\approx\tfrac{\nu_{i+1/2}}{2}\big{(}e_{i+1/2}^{-}-e_{i+1/2}^{+}\big{)}$ from (SM77), our next approximation is to replace the unknown points $\{\xi_{\ell}\}$ in (SM82) with a set of equidistant points such that we can discretize with finite differences the derivatives involved. For $k=1$, we simply set $\xi_{\ell}=x_{i+1/2}$. Recalling that the reconstruction weights are solution independent for $k=1$ with $b_{i}=\widetilde{b}_{i}=1$, plugging into (SM77) gives
$\widehat{E}_{i+1/2}\approx\frac{\nu_{i+1/2}}{2}\big{(}e_{i+1/2}^{-}-e_{i+1/2}^{+}\big{)}\approx-h\frac{\nu_{i+1/2}}{2}\frac{\operatorname{d}\!e}{\operatorname{d}\!x}\bigg{|}_{x_{i+1/2}}\quad{\rm for}\quad k=1.$ (SM83)
@@
Now consider $k>1$; in principle this could be done for arbitrary $k$, but since we only present numerical results for $k=2$, let us just consider that case. This case requires slightly more care because we have no reason to expect that $\xi_{0}$ and $\xi_{1}$ are a distance of $h$ apart. Introducing in (SM82) the ansatz that $\xi_{0}=\xi_{1}-h$ requires us to make the following adjustment:
@@
$e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $\approx\beta\frac{(-h)^{k}}{(k+1)!}\bigg{|}_{k=2}\sum\limits_{\ell=0}^{1}\Big{[}\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}-b_{i}^{\ell}\,\zeta^{\ell}\Big{]}\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg{|}_{\xi_{0}-\ell h}.$ (SM84)
@@
That is, we introduce the unknown coefficient $\beta$, which can be understood by appealing to the mean value theorem. To determine $\beta$, we enforce that when the optimal linear weights are used in (SM84), i.e., $b^{\ell}=d^{\ell},\widetilde{b}^{\ell}=\widetilde{d}^{\ell}$, that the result is the same as (SM59) when $k=2$. Recall that (SM59) estimates $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ when $e_{i+1/2}^{\pm}$ are weighted reconstructions using optimal linear weights. Setting the resulting two expressions equal after rearrangement gives
[…]
@@
Evidently, this is satisfied with $\beta=-\frac{3}{4}$ and $\xi_{0}=x_{i+1}$. Plugging the results into (SM84) and then (SM77) we arrive at:
$\widehat{E}_{i+1/2}\approx-\frac{3}{4}\frac{h^{2}}{6}\frac{\nu_{i+1/2}}{2}\bigg{[}\Big{(}2\widetilde{b}_{i+1}^{0}+b_{i}^{0}\Big{)}\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg{|}_{x_{i+1}}-\Big{(}\widetilde{b}_{i+1}^{1}+2b_{i}^{1}\Big{)}\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg{|}_{x_{i}}\bigg{]}\quad{\rm for}\quad k=2.$ (SM86)
@@
Recall from Footnote 10 that (^$\big{(}\widehat{{\cal E}}\bar{\bm{e}}\big{)}_{i}=\big{(}\widehat{E}_{i+1/2}-\widehat{E}_{i-1/2}\big{)}/h$. Thus, from (SM83) and (SM86) we have the following approximations
@@
$\sum\limits_{j=0}^{m-1}\widehat{{\cal E}}^{n+j}=\begin{cases}&\displaystyle{\frac{h}{2}}{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg{)}{\cal D}_{1}^{\top},\quad k=1,\\ \begin{split}&\frac{3}{4}\frac{h^{2}}{12}{\cal D}_{1}\bigg{[}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\odot\bm{\gamma}^{1,n+j}\bigg{)}{\cal D}_{2}^{1}\\ &\hskip 34.44434pt-\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\odot\bm{\gamma}^{0,n+j}\bigg{)}{\cal D}_{2}^{0}\bigg{]},\end{split}\quad k=2.\end{cases}$ (SM87)

### v2

@@ Acknowledgments
Figure SM7: Numerical test problems for MGRIT applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM7). Left column: The wave-speed function $\alpha(x,t)$. Right column: Space-time contour of the corresponding solution $e(x,t)$ of the PDE (cons-lin). From top to bottom, the wave-speed functions are: $\alpha(x)=1$, $\alpha(x)=\frac{1}{2}\big(1+\cos^{2}(\pi x)\big)$, $\alpha(x,t)=-\sin^{2}\big(\pi(x-t)\big)$, and $\alpha(x,t)=\cos(2\pi x)\cos(2\pi t)$, respectively.
In this section, we show results for MGRIT used to solve discretizations of the PDE (cons-lin) for four different wave-speed functions $\alpha$. The different wave-speed functions are shown in Fig. SM7, along with contours of the corresponding PDE solution in space-time. In all cases, the PDE initial condition is $e(x,0)=\sin^{4}(\pi x)$, the time domain is $t\in[0,4]$, and the (fine-grid) time-step is $\delta t=0.85h$.
@@
Figure SM8: Residual histories for MGRIT V-cycles applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM7). Left column: 1st-order accurate discretizations. Right column: 3rd-order accurate discretizations. Each of the four rows corresponds with the test problem shown in the corresponding row of Fig. SM7. Solid lines correspond to GLF numerical fluxes, and dashed lines correspond to LLF numerical fluxes. Note that for the constant wave-speed problem (top row), the GLF and LLF numerical fluxes are identical (see discussion in Section SM7.1).
@@
Figure SM9: Residual histories for MGRIT V-cycles applied to solve the linear conservation law (cons-lin) discretized with standard linear method-of-lines discretization (Section SM7). The test problems here are the same as the GLF problems considered in the first two rows of Fig. SM8, except that the coarse-grid operator does not include a truncation error correction. Note: All vertical axes are on the same scale ranging from $10^{-10}\to 10^{5}$, while those in Fig. SM8 range from $10^{-10}\to 1$.
MGRIT residual history plots are then shown in Fig. SM8. These results are for multilevel V-cycles, using a coarsening factor of $m=8$ on all levels, and coarsening is continued until fewer than two points in time would result. FCF-relaxation is used on all levels. Coarse-grid operators on deeper levels in the hierarchy have the same structure as the operator (SM74) on level two, with the truncation errors being computed recursively (see discussion in [14, Section 3.4]). All coarse-grid linear systems are approximately solved with GMRES, which is iterated until residual is reduced in norm by by 0.01 or until 10 iterations are reached. The parameter $\vartheta$ that specifies exactly where inside each time interval the wave-speed and dissipation coefficients is frozen is set to zero. The initial MGRIT iterate is a vector with uniformly random entries, and MGRIT is iterated until the 2-norm of the space-time residual is reduced by at least 10 orders of magnitude.
@@
For comparative purposes, we include in Fig. SM9 MGRIT residual histories for a subset of the same problems from Fig. SM8, except that no approximate truncation error correction has been used on coarse levels. That is, the coarse-grid operator used in the tests in Fig. SM9 is (SM74) with ${\cal T}_{\rm ideal}^{n}={\cal T}_{\rm direct}^{n}=0$, or simply the rediscretized semi-Lagrangian method, $\Psi^{n}={\cal S}^{n,m\delta t}_{p,1_{*}}$. Evidently, the approximate truncation error correction substantially improves the convergence rate of MGRIT, and appears necessary for the solver to converge in a number of iterations $k\ll n_{t}/(2m)$—the number of iterations at which FCF-relaxation sequentially propagates the initial condition across the entire time domain [20].
SM8 Approximating truncation error for linearized method-of-lines discretization
@@
In this section, we describe our approach for approximating the truncation error operator ${\cal T}_{\rm ideal}^{n}$ in the coarse-grid operator (40), in the context where the fine-grid operator is a linearized method-of-lines discretization. Recall that the particular details of the linearized problem are given in Section 4. This situation is distinct from that in Section SM7, which developed an approximation of ${\cal T}_{\rm ideal}^{n}$ in the context of the fine-grid operator being a standard, linear method-of-lines discretization. That being said, the heuristics we use in this section here are heavily inspired by the work in Section SM7 given the excellent results in Fig. SM8; we strongly recommend reading Section SM7 before the current section.
@@
As in Section SM7, we suppose that the approximate truncation error operator takes the form
(SM86) ${\cal T}_{\rm ideal}^{n}=\textrm{approximation}\bigg(\sum_{j=0}^{m-1}\Big(\delta t\widehat{{\cal E}}^{n+j}-\widehat{e}_{\rm RK}\big(\delta t\widehat{L}^{n+j}\big)^{p+1}\Big)\bigg).$
@@
Note that the truncation error operator in Section SM7 was written generally for a $q$th-order ERK method, but here we have fixed q=p=2k1q=p=2k-1 since we only consider discretizations of the nonlinear PDE (cons) with equal spatial and temporal orders of accuracy.
@@
The rest of this section is organized as follows. Section SM8.1 presents an approximation for spatial discretization errors $\widehat{{\cal E}}^{n+j}$ in (SM86). Section SM8.2 presents an approximation for the powers of the spatial discretization $\big(\widehat{L}^{n+j}\big)^{p+1}$ appearing in (SM86). The final approximation for (SM86) is presented in Section SM8.3.
SM8.1 Approximating the spatial discretization error
@@
Estimating rigorously the error term $\widehat{E}^{n+j}$ in the present setting is quite complicated. So, we instead pursue a heuristic approach based the result in Lemma SM7.3. Recall that Lemma SM7.3 estimates the error of the LF numerical flux (10) when: (i) reconstructions of the wave-speed in the flux are replaced with exact evaluations of the wave-speed, and (ii) the linear reconstructions $e_{i+1/2}^{\pm}$ use standard weights. This differs from the present setting, in which the LF numerical flux (10) uses: (i) reconstructions of the wave-speed that are based on the current linearization point (see Section 4.2), and (ii) the linear reconstructions $e_{i+1/2}^{\pm}$ use non-standard weights (see Section 4.3). Nonetheless, we suppose that the dominant error term in the numerical flux takes the same form as it does in the case of the standard linear discretization described above, and as is analyzed in Lemma SM7.3 of Section SM7. That is, we suppose (dropping temporal superscripts for readability): [^12] It is possible to analyze the error $\widehat{E}_{i+1/2}$ more rigorously than this; however, we do not present such details here since they do not appear necessary or helpful for our purposes.
(SM87) $\widehat{E}_{i+1/2}\approx\frac{\nu_{i+1/2}}{2}\big(e_{i+1/2}^{-}-e_{i+1/2}^{+}\big).$
@@
Thus, our first step for developing an expression for (SM86) is to estimate the difference $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ in (SM87). Recall from Section 4.3 that the exact form of the linearized reconstructions $e_{i+1/2}^{\pm}$ depends on the order of the reconstructions $u_{i+1/2}^{\pm}$ in the outer nonlinear iteration and on the linearization scheme being used. Specifically, if 1st-order reconstructions are used then $e_{i+1/2}^{\pm}$ are just standard linear reconstructions. For higher-order reconstructions $u_{i+1/2}^{\pm}$, the reconstructions are nonlinear since they depend on $\bar{\bm{u}}$ (supposing WENO weights are used, that is). The linearized reconstructions $e_{i+1/2}^{\pm}$ are then either: (i) approximate gradients of $u_{i+1/2}^{\pm}$ if the approximate Newton linearization (35) is used; or (ii) weighted reconstructions using the same weights as were used to compute $u_{i+1/2}^{\pm}$ if the Picard linearization (36) is used.
@@
(SM88) $e_{i+1/2}^{-}=\sum\limits_{\ell=0}^{k-1}b_{i}^{\ell}(\bar{\bm{u}})\big(R^{\ell}\bar{\bm{e}}\big)_{i},\quad e_{i+1/2}^{+}=\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\big(\widetilde{R}^{\ell}\bar{\bm{e}}\big)_{i+1}.$
Our starting point for estimating $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ is the following corollary of Lemma SM6.5, which considered estimates for the standard polynomial reconstructions $\big(\widetilde{R}^{\ell}\bar{\bm{e}}\big)_{i},\big({R}^{\ell}\bar{\bm{e}}\big)_{i}$.
Corollary SM8.1 (Error estimate for weighted reconstructions using 2k12k-1 cells).
@@
(SM89) $\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i}^{\ell}\big(\widetilde{R}^{\ell}\bar{\bm{e}}\big)_{i}=e(x_{i-1/2})-\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[\widetilde{b}_{i}^{\ell}\,\widetilde{\zeta}^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i,\ell}(x_{i-1/2})}\right],$
@@
(SM90) $\sum\limits_{\ell=0}^{k-1}{b}_{i}^{\ell}\big({R}^{\ell}\bar{\bm{e}}\big)_{i}=e(x_{i+1/2})-\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[b_{i}^{\ell}\,\zeta^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i,\ell}(x_{i+1/2})}\right].$
Proof SM8.2.
The result follows by substituting the error estimates of $\big(\widetilde{R}^{\ell}\bar{\bm{e}}\big)_{i}$ and $\big({R}^{\ell}\bar{\bm{e}}\big)_{i}$ given in Lemma SM6.5 and then using the consistency of the weights.
@@
(SM91) $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $=\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\left[\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i+1,\ell}(x_{i+1/2})}-b_{i}^{\ell}\,\zeta^{\ell}\,\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{i,\ell}(x_{i+1/2})}\right],$
(SM92) $\approx\frac{(-h)^{k}}{(k+1)!}\sum\limits_{\ell=0}^{k-1}\Big[\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}-b_{i}^{\ell}\,\zeta^{\ell}\Big]\frac{\operatorname{d}\!^{k}e}{\operatorname{d}\!x^{k}}\bigg|_{\xi_{\ell}},$
@@
where $\{\xi_{\ell}\}$ are unknown points. The approximation made here is that, for fixed $\ell$, we can replace the two unknown points $\xi_{i,\ell}(x_{i+1/2})$ and $\xi_{i+1,\ell}(x_{i+1/2})$ by single unknown point $\xi_{\ell}$. This is true to at least lowest order because $\xi_{i,\ell}(x_{i+1/2})$ and $\xi_{i+1,\ell}(x_{i+1/2})$ are at most a distance of ${\cal O}(h)$ apart (see Lemma SM6.5 for why this is true).
@@
Recalling that ultimately we want to discretize $\widehat{E}_{i+1/2}\approx\tfrac{\nu_{i+1/2}}{2}\big(e_{i+1/2}^{-}-e_{i+1/2}^{+}\big)$ from (SM87), our next approximation is to replace the unknown points $\{\xi_{\ell}\}$ in (SM92) with a set of equidistant points such that we can discretize with finite differences the derivatives involved. For $k=1$, we simply set $\xi_{\ell}=x_{i+1/2}$. Recalling that the reconstruction weights are solution independent for $k=1$ with $b_{i}=\widetilde{b}_{i}=1$, plugging into (SM87) gives
(SM93) $\widehat{E}_{i+1/2}\approx\frac{\nu_{i+1/2}}{2}\big(e_{i+1/2}^{-}-e_{i+1/2}^{+}\big)\approx-h\frac{\nu_{i+1/2}}{2}\frac{\operatorname{d}\!e}{\operatorname{d}\!x}\bigg|_{x_{i+1/2}}\quad{\rm for}\quad k=1.$
@@
Now consider $k>1$; in principle this could be done for arbitrary $k$, but since we only present numerical results for $k=2$, let us just consider that case. This case requires slightly more care because we have no reason to expect that $\xi_{0}$ and $\xi_{1}$ are a distance of $h$ apart. Introducing in (SM92) the ansatz that $\xi_{0}=\xi_{1}-h$ requires us to make the following adjustment:
@@
(SM94) $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ $\approx\beta\frac{(-h)^{k}}{(k+1)!}\bigg|_{k=2}\sum\limits_{\ell=0}^{1}\Big[\widetilde{b}_{i+1}^{\ell}\,\widetilde{\zeta}^{\ell}-b_{i}^{\ell}\,\zeta^{\ell}\Big]\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg|_{\xi_{0}-\ell h}.$
@@
That is, we introduce the unknown coefficient $\beta$, which can be understood by appealing to the mean value theorem. To determine $\beta$, we enforce that when the optimal linear weights are used in (SM94), i.e., $b^{\ell}=d^{\ell},\widetilde{b}^{\ell}=\widetilde{d}^{\ell}$, that the result is the same as (SM69) when $k=2$. Recall that (SM69) estimates $e_{i+1/2}^{-}-e_{i+1/2}^{+}$ when $e_{i+1/2}^{\pm}$ are weighted reconstructions using optimal linear weights. Setting the resulting two expressions equal after rearrangement gives
(SM95) $\frac{\operatorname{d}\!^{3}e}{\operatorname{d}\!x^{3}}\bigg|_{x_{i+1/2}}=-\beta\frac{1}{h}\frac{4}{3}\left[\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg|_{\xi_{0}}-\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg|_{\xi_{0}-h}\right]+\textrm{h.o.t.}$
@@
Evidently, this is satisfied with $\beta=-\frac{3}{4}$ and $\xi_{0}=x_{i+1}$. Plugging the results into (SM94) and then (SM87) we arrive at:
(SM96) $\widehat{E}_{i+1/2}\approx-\frac{3}{4}\frac{h^{2}}{6}\frac{\nu_{i+1/2}}{2}\bigg[\Big(2\widetilde{b}_{i+1}^{0}+b_{i}^{0}\Big)\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg|_{x_{i+1}}-\Big(\widetilde{b}_{i+1}^{1}+2b_{i}^{1}\Big)\frac{\operatorname{d}\!^{2}e}{\operatorname{d}\!x^{2}}\bigg|_{x_{i}}\bigg]\quad{\rm for}\quad k=2.$
@@
Recall from Footnote 10 that $\big(\widehat{{\cal E}}\bar{\bm{e}}\big)_{i}=\big(\widehat{E}_{i+1/2}-\widehat{E}_{i-1/2}\big)/h$. Thus, from (SM93) and (SM96) we have the following approximations
@@
(SM97) $\sum\limits_{j=0}^{m-1}\widehat{{\cal E}}^{n+j}=\begin{cases}&\displaystyle{\frac{h}{2}}{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\bigg){\cal D}_{1}^{\top},\quad k=1,\\ \begin{split}&\frac{3}{4}\frac{h^{2}}{12}{\cal D}_{1}\bigg[\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\odot\bm{\gamma}^{1,n+j}\bigg){\cal D}_{2}^{1}\\ &\hskip 34.44434pt-\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\nu}^{n+j}\odot\bm{\gamma}^{0,n+j}\bigg){\cal D}_{2}^{0}\bigg],\end{split}\quad k=2.\end{cases}$